# OMP4Py: a pure Python implementation of OpenMP

### v1

@@ 1 Introduction
A thorough experimental evaluation was carried out in the paper, demonstrating that OMP4Py has significant potential for hybrid parallelism in combination with mpi4py [4] and for non-numerical workloads. However, the current threading limitations of Python’s interpreter (v3.13) hinder its scalability for numerical applications.
@@ 2.1 OpenMP
OpenMP [3] is a parallel programming model designed for shared-memory computer systems, aiming to simplify the exploitation of inherent concurrency in many algorithms. OpenMP uses the fork-join model of parallel execution. In this model, the program starts with a single master thread. At parallel regions, the master thread creates multiple parallel threads that concurrently execute the assigned tasks. Once the parallel tasks are completed, the threads join back into the master thread, which continues executing the program sequentially. Through compiler directives, programmers can instruct the compiler to generate multithreaded code at a higher level. The OpenMP API is compatible with C/C++ and Fortran. Unlike lower-level approaches such as pthreads in the POSIX library, where the programmer manually manages thread creation and task assignment, OpenMP abstracts these complexities, providing a higher-level programming model. For this reason, OpenMP is considered by the HPC community as the standard for exploiting the multithreading capabilities of current multi-core CPUs.
@@
      #pragma omp parallel - Defines a parallel region, which is a block of code executed by multiple threads.
@@
      #pragma omp barrier - All threads must reach a barrier before any can continue.
@@
      firstprivate(list) - Similar to private, but the initial value of the variable is copied from the master thread.
@@
      shared(list) - Specifies that variables in the list should be shared among all threads.
@@
      reduction(op:list) - Performs a reduction operation on variables across all threads in a parallel region.
@@
      omp_set_num_threads(int) - Sets the number of threads that will be used in the next parallel region.
@@ 3 OMP4Py
OMP4Py is a novel implementation of the OpenMP standard, designed specifically for Python. It currently supports the complete specifications of version 3.0. This tool was developed with adherence to the official OpenMP documentation [^3] https://www.openmp.org/wp-content/uploads/spec30.pdf [online, accessed November 5, 2024], which outlines the necessary compiler directives, runtime library functions, and environment variables required for creating shared memory parallel programs using threads. The OpenMP standard traditionally supports C, C++, and Fortran, which are low-level, compiled languages that use preprocessor directives to manage parallelism. These directives are not part of the core language syntax but are preprocessor instructions that modify the source code prior to the compilation process. In C/C++, these directives are introduced using #pragma, and in Fortran, they are marked by the sentinel !\$. The main goal of OMP4py is to bring the familiar parallelization paradigm of OpenMP to Python, allowing Python developers to write parallel code with the same level of control and flexibility as in C, C++, or Fortran. This tool aims to port the OpenMP model, with all code execution handled natively using Python threads. This ensures integration with Python’s libraries, enabling multi-threaded performance directly within Python’s ecosystem.
@@ 3.1 Parallel Directive
In OpenMP, the parallel directive is essential for creating parallel programs. This construct allows a specific segment of code to execute simultaneously across multiple threads. When a code block is preceded by this construct, the program creates a team of threads, and each thread, including the original one, executes a copy of the code block concurrently. Multiple parallel directives can be nested, enabling each thread to create new teams of threads that work independently. This feature is known as nested parallelism and must be enabled with omp_set_nested. By definition in the standard, every OpenMP program runs within an implicit single-threaded parallel construct. This allows API functions like omp_get_num_threads or omp_get_thread_num, among others, to operate independently of where they are called in the code. If nested parallelism is disabled, only an explicit parallel execution is allowed; subsequent parallel regions are ignored and executed sequentially by the threads that encounter them.
Variables defined within a parallel block are local to each thread, meaning that each thread has its own copy. However, previously defined variables can be either private or shared, according to the user’s preference. Shared variables maintain the same value across all threads. Any modifications made by one thread are visible to the others, and this shared value persists even after the parallel block ends. By default, all variables defined before a parallel block are shared. While users can explicitly mark variables as shared, they only need to specify which variables should be private. Declaring a variable private makes it behave as if it was created inside the parallel block, causing its initial value to be lost when the block ends. To retain the value of a variable, users can employ firstprivate, which initializes each thread’s local copy to the variable’s previous value.
@@
The implementation of the parallel directive in Python presents two main challenges. First, the code that is meant to run in parallel across different threads must be placed inside a function. Second, it is necessary to manage variables used both inside and outside the code block to determine whether they should be shared or private for each thread. Figure 2 shows an example of a user-defined parallel block (top) and the corresponding code generated by OMP4Py for its parallel execution (bottom). It is important to note that, in code generation examples, the prefix _omp_ will be used for all internal OMP4Py symbols, with dynamically created symbols including a number to avoid collisions. The example shows a simple function with a parallel directive and six representative variables (’a’ to ’f’) with different types. The first change to observe is that the decorator (line 3, top code) and the function containing the directive (line 9, top code) have been removed. Once the code is transformed, the decorator is no longer needed and must be removed to prevent multiple processing. The parallel directive (line 9, top code) has been transformed into a function (line 8, bottom code) that will be called by different threads. The _omp_parallel_run function (line 18, bottom code) is an internal OMP4Py function responsible for initializing and launching threads using Python’s Thread class. Since the user has specified a number of threads with num_threads, we also need to pass this argument. Note that the existing thread, or master thread, is an execution thread as well, so only $num\_threads-1$ additional threads are created.
@@ 3.2 Worksharing Constructs
The distribution of work among threads is the base of OpenMP’s functionality. A worksharing construct divides execution regions among the team of threads created in the most recent parallel directive. Common worksharing constructs include for loops, which distribute loop iterations among threads, sections, which divide distinct code blocks among threads, and single, which ensures that a specific section of code is executed by only one thread. Worksharing regions are executed as soon as the first thread encounters them. While the first thread can begin executing its part of the work immediately, it must remain in the worksharing region until all threads in the team reach the same point. This is the default behavior for all worksharing constructs, although users can modify it if necessary using the nowait clause. If there is only a single thread, no work distribution occurs, and the worksharing regions are executed sequentially.
@@ 3.2.1 For
The first challenge to implement the for directive in Python is the absence of the traditional for loop. In Python, all for loops are constructed as foreach loops that iterate over a collection of elements. Consequently, to achieve the equivalent of a classic for loop, one must use a foreach loop over a range of values generated by the built-in range function. For example, a simple C-style loop like for(int i = 0; i < 10; i++) is equivalent to for i in range(0, 10, 1) in Python, where the initial value, final value, and step are specified as arguments.
@@
Figure 3 shows an example of a loop parallelized with the for directive (top) and its internal implementation generated by OMP4Py (bottom). First, we need to note that the for directive must be within a parallel directive to distribute the work (lines 6-7, top code). Unlike the parallel directive, the for directive must be the unique element within the with block; no other statements are allowed inside the block except within the loop itself. The implementation of this directive is simple: the range function (line 8, top code) is replaced by _omp_range (line 7, bottom code), and the scheduling and chunk values are passed as keyword arguments to the function. The _omp_range function is always generated with the three positional arguments of range, even if the user does not explicitly specify them. The function, implemented in the OMP4Py runtime, returns a different iterator for each thread, which behaves according to the selected scheduling policy by returning the values assigned to that thread one by one. Once the iterator runs out of elements, it blocks until all other iterators have finished. This synchronization is achieved because all threads share a barrier object: once an iterator completes, it waits at the barrier until all threads reach it. As previously mentioned, this behavior can be avoided using the nowait clause.
@@ 3.2.2 Sections
The sections directive divides the work among threads by assigning each thread a structured block defined with the section directive. Therefore, only blocks with the section directive can exist within a sections directive. Each block will be executed once by a single thread, but the order of execution is not predetermined, and any thread may execute one or multiple blocks depending on availability. In sequential execution, the blocks are executed in the order of their definition. Similar to the for directive, there is a synchronization barrier that blocks all threads until all section blocks have been executed. This barrier can also be removed using the nowait clause.
@@ 3.2.3 Single
Figure 6 shows an example of user code when using a single directive (top) and its internal implementation by OMP4Py (bottom). In this code, a variable declared with the firstprivate clause (line 6, top code) creates a private copy for each thread, initialized with the value from before the parallel region begins. Inside the single block, the variable is incremented by one (line 8, top code). After the single block finishes, the value of the variable from the thread that executed the block is broadcast to all other threads using the copyprivate clause. Note that variables listed in the copyprivate clause must be private to each thread. Once the broadcast is complete, all threads have the updated value and print it (line 9, top code).
@@ 3.3 Tasking Directives
Figure 7 shows the recursive calculation of Fibonacci numbers using tasks (top) and its internal implementation generated by OMP4Py (bottom). The algorithm defines the Fibonacci function (line 4, top code), which calculates the value for a given $n$ by summing the recursive calls to the same function for the values $n-1$ and $n-2$ (lines 10 and 12, top code). The initial call to the Fibonacci function is placed in a single block to ensure it is executed by only one thread (line 21, top code), while the remainder threads are blocked. The function then creates a task for the recursive calculation (lines 9 and 11, top code) and blocks the current task until they have been executed (line 13, top code). The threads in the team will consume the created tasks and generate new tasks until they reach the base cases of $n=0$ and $n=1$. The code generated by OMP4Py shows the similarities with the parallel directive. In this way, the code within the task directive (lines 9 and 11, top code) and the parallel directive (line 19, top code) both result in the creation of functions: _omp_task_1, _omp_task_2 (lines 8 and 13 in the bottom code), and _omp_parallel_1 (line 23, bottom code), respectively. The only notable difference is that parallel uses the function _omp_parallel_run (line 29, bottom code), while task employs _omp_task_submit (lines 12 and 17, bottom code) which places the task function into a shared queue. The queue is accessible by all threads that will be able to pick up the task and execute it. Finally, the taskwait clause is implemented as a simple function (line 18, bottom code), which forces the caller thread to consume all tasks in the queue until it is empty, at which point the function returns.
@@ 3.4 Interaction with Python
When a Python program using OMP4Py is executed, a single thread begins by running the main program, as in a typical Python script. However, it is important to note that before the main program starts, the Python interpreter processes all decorators, including omp decorators. So, by the time the main program starts, all OpenMP directives within the source code will be replaced with a new pure Python version with parallel behavior. According to the OpenMP standard, this initial thread is called the master thread. OMP4Py initializes the master thread context with the first call to any function in its API. The remaining threads created in parallel constructs will be initialized with a context derived from the master thread before they begin execution. For this reason, any thread created outside the OMP4Py clauses, for instance, using asyncio, concurrent.futures, threading, or the multiprocessing module, will not have an associated context and will be treated by OMP4Py as an uninitialized master thread. If there is a subsequent call to any function in the OMP4Py API by one of these threads, a new OMP4Py context will be created for it. In this way, the thread will act as a new master thread and can independently use all the features of OMP4Py, leaving the user responsible for managing potential concurrency issues between different instances.
@@
The context in OMP4Py is implemented by a stack object where OpenMP construct tasks are stored. In the OpenMP standard, the master thread exists within a single-threaded parallel region. Therefore, the initialization of the master thread’s context involves adding a parallel construct task to the stack, allowing the OpenMP API (like omp_get_thread_num) to be called from any point in the source code, ensuring consistent results. The context is stored locally to the thread using threading.local.
@@
In a typical execution flow of an OpenMP application, the next task added to the stack is the parallel task, which creates a parallel region with multiple threads. The _omp_parallel_run function is responsible for this process. It works as follows: the master thread creates a context for each new thread, with a stack that is a shadow copy of its own. The master thread then adds the parallel task to all stacks, including its own, and finally, using the threading library, the threads are created with their respective contexts as arguments. The parallel task contains the following information: the thread ID, the number of threads, a mutex, a barrier, a shared task list, and a shared dictionary. The thread ID and number of threads are critical for work distribution and API functions. The mutex and barrier are used for constructs like critical and barrier. The shared task list holds tasks (created by the task construct) that must be completed before the parallel region ends, while the shared dictionary stores common data, such as iteration counters for dynamic workloads in a for construct. Exceptions thrown within a parallel region are not automatically propagated to the master thread. Instead, the behavior depends on how the exception is handled within each thread: if an exception occurs, it must be caught and handled within that thread. OMP4Py catches unhandled exceptions in parallel regions to prevent the program from terminating unexpectedly, but relying on this behavior is considered bad practice.

### v2

@@ 1 Introduction
A thorough experimental evaluation was carried out in the paper, demonstrating that OMP4Py has significant potential for hybrid parallelism in combination with mpi4py [4] and for non-numerical workloads. However, the current threading limitations of Python’s interpreter (v3.13) hinder its scalability for numerical applications. It is important to highlight that as these limitations in the Python interpreter are progressively resolved, the scalability constraints of OMP4Py when running numerical applications will gradually disappear, without requiring any modifications to its implementation.
@@ 2.1 OpenMP
OpenMP [3] is a parallel programming model originally designed for shared-memory computer systems, aiming to simplify the exploitation of inherent concurrency in many algorithms. OpenMP primarily follows the fork-join model of parallel execution. In this model, the program starts with a single initial thread. At parallel regions, the initial thread creates multiple parallel threads that concurrently execute the assigned tasks. Once the parallel tasks are completed, the threads join back into the initial thread, which continues executing the program sequentially. Through compiler directives, programmers can instruct the compiler to generate multithreaded code at a higher level of abstraction, avoiding the manual management of thread creation and task assignment required by low-level approaches like pthreads in the POSIX library. The OpenMP API is compatible with C/C++ and Fortran.
Although initially designed for shared-memory architectures, OpenMP has expanded its capabilities to support heterogeneous computing. Since the introduction of the target directive family, OpenMP has enabled offloading computations to accelerators, such as GPUs, which often employ distributed-memory models internally. This extension allows OpenMP to be used in hybrid computing environments where both shared- and distributed-memory paradigms coexist. For this reason, OpenMP remains the standard for exploiting multithreading capabilities of modern multi-core CPUs while also enabling high-performance execution on heterogeneous architectures.
@@
      #pragma omp parallel - Defines a parallel region, which is a block of code executed by a team of multiple threads. According to OpenMP terminology, a primary thread is the thread that encounters a parallel construct, creates a team of threads, generates a set of implicit tasks, and executes one of those tasks as thread number 0. If there is only a single team of threads, the initial thread and the primary thread refer to the same thread.
@@
      #pragma omp barrier - All threads in the current team must reach a barrier before any can continue.
@@
      firstprivate(list) - Similar to private, but the variable is initialized using the value from the initial thread.
@@
      shared(list) - Specifies that variables in the list should be shared among all threads in the current team.
@@
      reduction(op:list) - Performs a reduction operation on variables across all threads in a team in a parallel region.
@@
      omp_set_num_threads(int) - Sets the number of threads to be used in subsequent parallel regions, unless overridden by a more specific mechanism (e.g., the num_threads clause in a parallel directive) or modified by another call to omp_set_num_threads.
@@ 3 OMP4Py
OMP4Py is a novel implementation of the OpenMP standard, designed specifically for Python. It currently supports the complete specifications of version 3.0. As explained in Section 2.1, this ensures that we cover most of the needs of OpenMP HPC applications and programmers. This tool was developed with adherence to the official OpenMP documentation [^3] https://www.openmp.org/wp-content/uploads/spec30.pdf [online, accessed November 5, 2024], which outlines the necessary compiler directives, runtime library functions, and environment variables required for creating shared memory parallel programs using threads. The OpenMP standard traditionally supports C, C++, and Fortran, which are low-level, compiled languages that use specific syntax for parallelism directives, as defined in the language standard. These directives are part of the OpenMP specification, and while they may differ in syntax between languages (such as #pragma in C/C++ or sentinel !\$ in Fortran), they all serve the same purpose: to manage parallelism during compilation. The main goal of OMP4py is to bring the familiar parallelization paradigm of OpenMP to Python, allowing Python developers to write parallel code with the same level of control and flexibility as in C, C++, or Fortran. This tool aims to port the OpenMP model, with all code execution handled natively using Python threads. This ensures integration with Python’s libraries, enabling multi-threaded performance directly within Python’s ecosystem.
@@ 3.1 Parallel Directive
In OpenMP, the parallel directive is essential for creating parallel programs. This construct allows a specific segment of code to execute simultaneously across multiple threads. When a code block is preceded by this construct, the program creates a team of threads, and each thread, including the original one, executes a copy of the code block concurrently. Multiple parallel directives can be nested, enabling each thread to create new teams of threads that work independently. This feature is known as nested parallelism and must be enabled with omp_set_nested. By definition in the standard, every OpenMP program runs within an implicit single-threaded parallel construct. This allows API functions like omp_get_num_threads or omp_get_thread_num, among others, to operate independently of where they are called in the code. If multiple nested parallel directives are present and omp_set_nested is disabled, only the outermost directive spawns threads. Subsequent parallel regions do not create new threads; instead, only the existing threads that encounter the directive will execute the code.
Variables defined within a parallel block are local to each thread, meaning that each thread has its own copy. However, previously defined variables can be either private or shared, according to the user’s preference. Shared variables maintain the same value across all threads in the current team. Any modifications made by one thread are visible to the others, and this shared value persists even after the parallel block ends. By default, all variables defined before a parallel block are shared. While users can explicitly mark variables as shared, they only need to specify which variables should be private. Declaring a variable private makes it behave as if it were created inside the parallel block, with each thread starting with an uninitialized value. Within the parallel region, the private variable is treated independently by each thread, and its final value is discarded after the block ends. The outer variable remains unaffected by any changes made to the private copy during the execution. To retain the value of the outer variable, users can employ firstprivate, which initializes each thread’s local copy to the variable’s previous value.
@@
The implementation of the parallel directive in Python presents two main challenges. First, the code that is meant to run in parallel across different threads must be placed inside a function. Second, it is necessary to manage variables used both inside and outside the code block to determine whether they should be shared or private for each thread. Figure 2 shows an example of a user-defined parallel block (top) and the corresponding code generated by OMP4Py for its parallel execution (bottom). It is important to note that, in code generation examples, the prefix _omp_ will be used for all internal OMP4Py symbols, with dynamically created symbols including a number to avoid collisions. The example shows a simple function with a parallel directive and six representative variables (’a’ to ’f’) with different types. The first change to observe is that the decorator (line 3, top code) and the function containing the directive (line 9, top code) have been removed. Once the code is transformed, the decorator is no longer needed and must be removed to prevent multiple processing. The parallel directive (line 9, top code) has been transformed into a function (line 8, bottom code) that will be called by different threads. The _omp_parallel_run function (line 18, bottom code) is an internal OMP4Py function responsible for initializing and launching threads using Python’s Thread class. Since the user has specified a number of threads with num_threads, we also need to pass this argument. Note that the existing thread, or initial thread, is an execution thread as well, so only $num\_threads-1$ additional threads are created.
@@ 3.2 Worksharing Constructs
The distribution of work among threads is the base of OpenMP’s functionality. A worksharing construct divides execution regions among the team of threads created in the most recent parallel directive. Common worksharing constructs include for loops, which distribute loop iterations among threads, sections, which divide distinct code blocks among threads, and single, which ensures that a specific section of code is executed by only one thread. Worksharing regions are executed as soon as the first thread encounters them. While the first thread can begin executing its part of the work immediately, it must remain in the worksharing region until all threads in the team reach the same point. This is the default behavior for all worksharing constructs, although users can modify it if necessary using the nowait clause. If there is only a single thread, the work is still distributed according to the scheduling policy, although in this case, the work will be executed sequentially since only one thread is available to process it.
@@ 3.2.1 For
The first challenge to implement the for directive in Python is the absence of the traditional for loop. In Python, all for loops are constructed as foreach loops that iterate over a collection of elements. Consequently, to achieve the equivalent of a classic for loop, one must use a foreach loop over a range of values generated by the built-in range function. For example, a simple C-style loop like for(int i = 0; i < 10; i++) is equivalent to for i in range(0, 10, 1) in Python, where the initial value, final value, and step are specified as arguments. It is important to note that ranges cannot be used directly to divide the work because Python iterators are sequential, while C++ iterators allow random access, enabling direct access to any element, size calculation, and arithmetic operations between iterators.
@@
Figure 3 shows an example of a loop parallelized with the for directive (top) and its internal implementation generated by OMP4Py (bottom). First, we need to note that the for directive must be within a parallel directive to distribute the work (lines 6-7, top code). Unlike the parallel directive, the for directive must be the unique element within the with block; no other statements are allowed inside the block except within the loop itself. The implementation of this directive is simple: the range function (line 8, top code) is replaced by _omp_range (line 7, bottom code), and the scheduling and chunk values are passed as keyword arguments to the function. The _omp_range function is always generated with the three positional arguments of range, even if the user does not explicitly specify them. The function, implemented in the OMP4Py runtime, returns a different iterator for each thread, which behaves according to the selected scheduling policy by returning the values assigned to that thread one by one. Once the iterator runs out of elements, it blocks until all other iterators have finished. This synchronization is achieved because all threads in the current team share a barrier object: once an iterator completes, it waits at the barrier until all threads reach it. As previously mentioned, this behavior can be avoided using the nowait clause.
@@ 3.2.2 Sections
The sections directive divides the work among threads by assigning each thread a structured block defined with the section directive. Therefore, only blocks with the section directive can exist within a sections directive. Each block will be executed once by a single thread, but the order of execution is not predetermined, and any thread may execute one or multiple blocks depending on availability. In sequential execution, the blocks are executed in the order of their definition. Similar to the for directive, there is a synchronization barrier that blocks all threads in the current team until all section blocks have been executed. This barrier can also be removed using the nowait clause.
@@ 3.2.3 Single
Figure 6 shows an example of user code when using a single directive (top) and its internal implementation by OMP4Py (bottom). In this code, a variable declared with the firstprivate clause (line 6, top code) creates a private copy for each thread, initialized with the value from before the parallel region begins. Inside the single block, the variable is incremented by one (line 8, top code). After the single block finishes, the value of the variable from the thread that executed the block is broadcast to all other threads in the team using the copyprivate clause. Note that variables listed in the copyprivate clause must be private to each thread. Once the broadcast is complete, threads in the team have the updated value and print it (line 9, top code).
@@ 3.3 Tasking Directives
Figure 7 shows the recursive calculation of Fibonacci numbers using tasks (top) and its internal implementation generated by OMP4Py (bottom). The algorithm defines the Fibonacci function (line 4, top code), which calculates the value for a given $n$ by summing the recursive calls to the same function for the values $n-1$ and $n-2$ (lines 10 and 12, top code). The initial call to the Fibonacci function is placed in a single block to ensure it is executed by only one thread (line 21, top code), while the remainder threads are blocked. The function then creates a task for the recursive calculation (lines 9 and 11, top code) and blocks the current task until they have been executed (line 13, top code). The threads in the team will consume the created tasks and generate new tasks until they reach the base cases of $n=0$ and $n=1$. The code generated by OMP4Py shows the similarities with the parallel directive. In this way, the code within the task directive (lines 9 and 11, top code) and the parallel directive (line 19, top code) both result in the creation of functions: _omp_task_1, _omp_task_2 (lines 8 and 13 in the bottom code), and _omp_parallel_1 (line 23, bottom code), respectively. The only notable difference is that parallel uses the function _omp_parallel_run (line 29, bottom code), while task employs _omp_task_submit (lines 12 and 17, bottom code) which places the task function into a shared queue. The queue is accessible by all threads in the team that will be able to pick up the task and execute it. Finally, the taskwait clause is implemented as a simple function (line 18, bottom code), which forces the caller thread to consume all tasks in the queue until it is empty, at which point the function returns.
@@ 3.4 Interaction with Python
When a Python program using OMP4Py is executed, a single thread begins by running the main program, as in a typical Python script. However, it is important to note that before the main program starts, the Python interpreter processes all decorators, including omp decorators. So, by the time the main program starts, all OpenMP directives within the source code will be replaced with a new pure Python version with parallel behavior. According to the OpenMP standard, this thread is called the initial thread. OMP4Py initializes the initial thread context with the first call to any function in its API. The remaining threads created in parallel constructs will be initialized with a context derived from the initial thread before they begin execution. For this reason, any thread created outside the OMP4Py clauses, for instance, using asyncio, concurrent.futures, threading, or the multiprocessing module, will not have an associated context and will be treated by OMP4Py as an uninitialized initial thread. If there is a subsequent call to any function in the OMP4Py API by one of these threads, a new OMP4Py context will be created for it. In this way, the thread will act as a new initial thread and can independently use all the features of OMP4Py, leaving the user responsible for managing potential concurrency issues between different instances.
@@
The context in OMP4Py is implemented by a stack object where OpenMP construct tasks are stored. In the OpenMP standard, the initial thread exists within a single-threaded parallel region. Therefore, the initialization of the initial thread’s context involves adding a parallel construct task to the stack, allowing the OpenMP API (like omp_get_thread_num) to be called from any point in the source code, ensuring consistent results. The context is stored locally to the thread using threading.local.
@@
In a typical execution flow of an OpenMP application, the next task added to the stack is the parallel task, which creates a parallel region with multiple threads. The _omp_parallel_run function is responsible for this process. It works as follows: the initial thread creates a context for each new thread, with a stack that is a shadow copy of its own. The initial thread then adds the parallel task to all stacks, including its own, and finally, using the threading library, the threads are created with their respective contexts as arguments. The parallel task contains the following information: the thread ID, the number of threads, a mutex, a barrier, a shared task list, and a shared dictionary. The thread ID and number of threads are critical for work distribution and API functions. The mutex and barrier are used for constructs like critical and barrier. The shared task list holds tasks (created by the task construct) that must be completed before the parallel region ends, while the shared dictionary stores common data, such as iteration counters for dynamic workloads in a for construct. Exceptions thrown within a parallel region are not automatically propagated to the initial thread. Instead, the behavior depends on how the exception is handled within each thread: if an exception occurs, it must be caught and handled within that thread. OMP4Py catches unhandled exceptions in parallel regions to prevent the program from terminating unexpectedly, but relying on this behavior is considered bad practice.